OpenStudy (maxdrizner):

A system of equations is shown below: x + 3y = 5 (equation 1) 7x - 8y = 6 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? (4 points)

OpenStudy (maxdrizner):

Show that the solution to the system of equations 3x + y = 5 and 8x -7y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 8x - 5y = 11 and 7x - 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations 15x + 13y = 17 and 7x - 8y = 6 is the same as the solution to the given system of equations Show that the solution to the system of equations -13x + 15y = 17 and 7x - 8y = 6 is the same as the solution to the given system of equations

OpenStudy (lena772):

Do you remember the answer?

OpenStudy (lena772):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

( ._____________________________.) lena U_U

OpenStudy (the_fizicx99):

He hasn't been on in a month, for all I know he could have gotten hit by bus >.<

OpenStudy (lena772):

i think he(ME) still needs assistance